# reference.bib

### v2

@@ A.2 Bump-on-Tail example
Figure 33: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with near initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 35: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with local initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 37: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EE) with far initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 38: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EE) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 39: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EE) with near initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 40: Simulation of (2.1) using (EE) with under-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 41: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EE) with local initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$,$H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 42: Simulation of (2.1) using (EE) with under-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 43: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with far initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 45: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with near initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 47: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with local initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 49: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EE) with far initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(T,x)$, $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$ and, convergence of objective.
@@
Figure 50: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EE) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 51: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EE) with near initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 52: Simulation of (2.1) using (EE) with over-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 53: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EE) with local initialization using GD with linesearch. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 54: Simulation of (2.1) using (EE) with over-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).

### v3

@@ B.2 Bump-on-Tail example
Figure 48: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 50: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (KL) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 52: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 53: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 54: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 55: Simulation of (2.1) using ($E$) with under-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 56: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$,$H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 57: Simulation of (2.1) using ($E$) with under-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 58: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 59: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 60: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 61: Simulation of (2.1) using (EET) with under-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 62: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$,$H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 63: Simulation of (2.1) using (EET) with under-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 64: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 66: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 68: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (KL) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 70: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using ($E$) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(T,x)$, $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$ and, convergence of objective.
@@
Figure 71: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using ($E$) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 72: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using ($E$) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 73: Simulation of (2.1) using ($E$) with over-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 74: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using ($E$) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
@@
Figure 75: Simulation of (2.1) using ($E$) with over-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 76: Simulation of (2.1) with under-parametrized $H$ obtained from (2.6) using (EET) with far initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(T,x)$, $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$ and, convergence of objective.
Figure 77: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EET) with far initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 78: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EET) with near initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 79: Simulation of (2.1) using (EET) with over-parametrized $H$ obtained from (2.6) with near initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 80: Simulation of (2.1) with over-parametrized $H$ obtained from (2.6) using (EET) with local initialization using GD with line-search. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).
Figure 81: Simulation of (2.1) using (EET) with over-parametrized $H$ obtained from (2.6) with local initialization using GD with constant stepsize. From left to right and top to bottom: $f[H](T=30,x,v)$, $|f[H](T,x,v)-f_{\text{eq}}(v)|$, $H$ and $E_{f[H]}(t,x)$, $\mathcal{E}_{f[H]}(t)$, convergence of objective and, trajectory over the landscape of the objective (yellow dot is initial guess).